# Boosting Diffusion-Based Text Image Super-Resolution Model Towards Generalized Real-World Scenarios

### v2

@@ 3.4 Loss Function
Table 1: Quantitative comparison for the synthetic dataset on CTR-TSR-Test [42] with different methods.
4 Experiments
4.1 Experimental Setup
Datasets In this work, we mainly focus on the SR of text characters in real-world scenarios as the exploration for complex characters, including Chinese characters. Following [47], we preprocess the image datasets with the following steps: (1) remove the images with a resolution smaller than 64 pixels; (2) only retain the images with a length of texts not larger than 24; (3) only retain the images with a width-to-height greater than 2; (4) resize the image to $128\times 512$. Then, we have over 400K HR text images, and each of these images has a paired LR image. The degradation to generate $\hat{x_{0}}$ and $\hat{y}$ is proposed in BSRGAN [44]. We divide the dataset into training and testing sets, with 80% of the instances divided in the training set and the remaining instances divided in the testing set. We also include the CTR-TSR-Test dataset from [42] for evaluation.
Evaluation Metrics and Baselines. To assess the performance of our proposed model, we compare it with several state-of-the-art image SR methods, including general SR approaches (e.g., ESRGAN [36], NAFNet [5]) and recent text image SR techniques (e.g., SRCNN [6], TSRN [35], TSBRN [4], MACRONet [21], DiffTSR [47]). For the baselines without training code, we adopt their released models; for these with training codes, we finetune them with our collected dataset for a fair comparison.
We utilize five evaluation metrics to quantify the effectiveness of these methods. First, we compute the peak signal-to-noise ratio (PSNR) and the learned perceptual image patch similarity (LPIPS) [45]. To further assess the realism of the generated images, we report the Frechet Inception Distance (FID) [10]. For a more specific evaluation of the text fidelity in restored text images, we incorporate word accuracy (ACC) and normalized edit distance (NED) [27]. Notably, we use the pre-trained TransOCR [4, 42] as the text recognition model to compute ACC and NED.
Implementation Details. Our method is implemented using PyTorch. All experiments are conducted on a multi-GPU setup with four NVIDIA A100 GPUs, each equipped with 80GB of memory. We set the batch size to 64 and use the Adam optimizer [18] with a learning rate of $1\times 10^{-5}$. In the loss function, the hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are empirically set to 1, 1, and 0.02, respectively.
@@ 4.2 Quantitative Comparison
Figure 4: Qualitative comparison on the real-world dataset with different methods including SRCNN [6], NAFNet [5], ESRGAN [36], the existing SOTA method DiffTSR [47] and our method for ×$\times 4$ super-resolution.
@@ 4.3 Qualitative Comparison
The qualitative results on the real-world dataset are shown in Figure 4. Our method outperforms all other SR and text super-resolution techniques, including SRCNN [6], NAFNet [5], ESRGAN [36], and the SOTA method DiffTSR [47], in generating high-quality text images. While ESRGAN [36] benefits from GAN-based generation to restore more visually realistic images, some results still exhibit noticeable artifacts, as seen in the second and third results in the third row. Similarly, DiffTSR [47] struggles to preserve realistic image qualities and often generates erroneous text structures.
In contrast, our method, with its strong SR prior and cross-attention mechanism, effectively restores text images with both high style realness and text fidelity. By generalizing well to real-world scenarios, our model produces more accurate and coherent results, surpassing the limitations of existing methods in handling real-world degradation.
@@ 4.4 Ablation Study
Remove Cross-attention Module. For the ‘w/o cross-attention’ experiment where we remove the cross-attention module and perform image recovery using only image features.
As illustrated in the second row of Figure 6, the characters are not well-formed and do not adhere to the correct structure of Chinese characters, leading to visually unpleasant results. This is because the model does not incorporate textual priors during the diffusion process, resulting in distorted character shapes and a lack of fidelity in the text.
@@
Without Progressive Sampling. In the ‘w/o progressive sampling’ experiment where we directly use the LR images to train the model, the model is hard to converge, which can be found in the fourth row of Figure 6. Directly training with LR images alone might not provide the model with enough difference between input and ground truth, making it harder for the model to learn fine-grained features necessary for high-quality image restoration.
@@
Also, the quantitative results are displayed in Table 2. We can see that by incorporating these components, our method is able to better utilize textual features, leading to more realistic and accurate text restoration. This demonstrates the importance of these three parts.
@@
w/o cross-attention 19.54 0.288 25.26 0.7105 0.6934
w/o confidence score 21.83 0.201 7.150 0.7047 0.7136
w/o progressive sampling 19.31 0.345 20.22 0.7415 0.7023
@@ Appendix A Training Details
Basic model information. All the models are trained on the real-world dataset. LR and HR images are resized into $128\times 512\times 3$ and $32\times 128\times 3$, respectively. Besides, we utilize the TDM and MoM module from [47] and load their pre-trained weights to process the text information and the text diffusion. When training the model, these two modules are frozen. For the Unet module, we utilize the pre-trained weights from the ResShift [43] model. Also, to fuse the textual feature from the MoM module with the image feature, after each basic layers on the ResShift, we insert the Transformer layer for the cross attention module. The Transformer layer is randomly initialized. During training, only the Unet module is updated. Besides, in the modeling of the text sequence, the maximum length of the text is set to 24 and all the characters in the text sequence belong to an alphabet with $K=6736$ characters including both Chinese and English characters as well as the numbers and special characters. The total step $T$ is set to 18.
@@ Appendix B Procedure of Our Framework
5:     Generate $\hat{x_{0}}$ and $\hat{y}$ via the pipeline in ESRGAN [36].
@@ Appendix D Future Works
Due to the limited performance of the OCR module, existing methods still make mistakes in the generated images. Therefore, in future work, we plan to incorporate reinforcement learning techniques to refine and correct the generated text, such as DPO [2]. Additionally, we aim to replace the OCR module with outputs from multi-modal large language models (MLLMs) [23, 24], which could potentially yield more accurate text outputs from the LR images.
@@ Appendix E More Visual Results
Figure 7: Qualitative comparison on the real-world dataset with different methods including SRCNN [6], NAFNet [5], ESRGAN [36], the existing SOTA method DiffTSR [47] and our method for ×$\times 4$ super-resolution.
Figure 8: Qualitative comparison on the real-world dataset with different methods including SRCNN [6], NAFNet [5], ESRGAN [36], the existing SOTA method DiffTSR [47] and our method for ×$\times 4$ super-resolution.
Figure 9: Qualitative comparison on the synthetic dataset with different methods including SRCNN [6], NAFNet [5], ESRGAN [36], the existing SOTA method DiffTSR [47] and our method for ×$\times 4$ super-resolution.

### v4

@@ 4.2 Quantitative Comparison
Figure 4: Qualitative comparison on the real-world dataset with different methods including SRCNN Dong et al. (2015), NAFNet Chen et al. (2022), ESRGAN Wang et al. (2021b), the existing SOTA method DiffTSR Zhang et al. (2024b) and our method for ×$\times 4$ super-resolution.
@@ 4.3 Qualitative Comparison
The qualitative results on the real-world dataset are shown in Figure 4. Our method outperforms all other SR and text super-resolution techniques, including SRCNN Dong et al. (2015), NAFNet Chen et al. (2022), ESRGAN Wang et al. (2021b), and the SOTA method DiffTSR Zhang et al. (2024b), in generating high-quality text images. While ESRGAN Wang et al. (2021b) benefits from GAN-based generation to restore more visually realistic images, some results still exhibit noticeable artifacts, as seen in the second and third results in the third row. Similarly, DiffTSR Zhang et al. (2024b) struggles to preserve realistic image qualities and often generates erroneous text structures. In contrast, our method, with its strong SR prior and cross-attention mechanism, effectively restores text images with both high style realness and text fidelity. By generalizing well to real-world scenarios, our model produces more accurate and coherent results, surpassing the limitations of existing methods in handling real-world degradation.
To provide additional qualitative evaluation, we conducted a human assessment study on a randomly selected subset of 100 images from our test dataset. The perceptual evaluation results demonstrate that our method generates images with acceptable quality in 70% of cases, while DiffTSR achieves this standard in only 25% of instances. This improvement in perceptual quality acceptance rate highlights the superior performance of our approach in generating visually plausible results.
@@ 4.4 Ablation Study
Remove Cross-attention Module. For the ‘w/o cross-attention’ experiment where we remove the cross-attention module and perform image recovery using only image features. As illustrated in the second row of Figure 5, the characters are not well-formed and do not adhere to the correct structure of Chinese characters, leading to visually unpleasant results. This is because the model does not incorporate textual priors during the diffusion process, resulting in distorted character shapes and a lack of fidelity in the text.
@@
Without Progressive Sampling. In the ‘w/o progressive sampling’ experiment, where LR images are directly employed to train the model, convergence proves to be challenging, as illustrated in the fourth row of Figure 5. Training exclusively with LR images may not sufficiently expose the model to the contrasts between input and ground truth, thereby hindering its ability to learn the intricate features essential for high-quality image restoration.
@@
Also, the quantitative results on CTR-TSR-Test are displayed in Table 2. We can see that by leverging these components, our method is able to better utilize textual features, leading to more realistic and accurate text restoration. Although our model utilizes a lightweight Unet module, it is still able to achieve satisfactory results after training. This demonstrates the importance of the three parts.
@@
w/o cross-attention 21.54 0.250 25.26 0.7105 0.6934
w/o confidence score 21.83 0.185 7.150 0.7047 0.7136
w/o progressive sampling 19.31 0.245 20.22 0.7415 0.7023
@@ Appendix A Training Details
Basic model information. All the models are trained on the real-world dataset. LR and HR images are resized into $128\times 512\times 3$ and $32\times 128\times 3$, respectively. Besides, we utilize the TDM and MoM module from Zhang et al. (2024b) and load their pre-trained weights to process the text information and the text diffusion. When training the model, these two modules are frozen. For the Unet module, we utilize the pre-trained weights from the ResShift Yue et al. (2024) model. Also, to fuse the textual feature from the MoM module with the image feature, after each basic layers on the ResShift, we insert the Transformer layer for the cross attention module. The inserted Transformer layer is randomly initialized. During training, only the Unet module and the MoM module is updated. Besides, in the modeling of the text sequence, the maximum length of the text is set to 24 and all the characters in the text sequence belong to an alphabet with $K=6736$ characters including both Chinese and English characters as well as the numbers and special characters. The total step $T$ is set to 18.
@@ Appendix B Procedure of Our Framework
5:     Generate $\hat{x_{0}}$ and $\hat{y}$ via the pipeline in ESRGAN Wang et al. (2021b).
@@ Appendix D Future Works
Due to the limited performance of the OCR module, existing methods still make mistakes in the generated images. Therefore, in future work, we plan to incorporate reinforcement learning techniques to refine and correct the generated text, such as DPO Black et al. (2023). Additionally, we aim to replace the OCR module with outputs from multi-modal large language models (MLLMs) Liu et al. (2023); Lu et al. (2024), which could potentially yield more accurate text outputs from the LR images.
@@ Appendix E More Visual Results
Figure 7: Qualitative comparison on the real-world dataset with different methods including SRCNN Dong et al. (2015), NAFNet Chen et al. (2022), ESRGAN Wang et al. (2021b), the existing SOTA method DiffTSR Zhang et al. (2024b) and our method for ×$\times 4$ super-resolution.
Figure 8: Qualitative comparison on the real-world dataset with different methods including SRCNN Dong et al. (2015), NAFNet Chen et al. (2022), ESRGAN Wang et al. (2021b), the existing SOTA method DiffTSR Zhang et al. (2024b) and our method for ×$\times 4$ super-resolution.
Figure 9: Qualitative comparison on the synthetic dataset with different methods including SRCNN Dong et al. (2015), NAFNet Chen et al. (2022), ESRGAN Wang et al. (2021b), the existing SOTA method DiffTSR Zhang et al. (2024b) and our method for ×$\times 4$ super-resolution.